### v4

# Causal Graph Neural Networks for Healthcare

Abstract
Healthcare artificial intelligence systems routinely fail when deployed across institutions, with documented performance drops and perpetuation of discriminatory patterns embedded in historical data. This brittleness stems, in part, from learning statistical associations rather than causal mechanisms. Causal graph neural networks address this triple crisis of distribution shift, discrimination, and inscrutability by combining graph-based representations of biomedical data with causal inference principles to learn invariant mechanisms rather than spurious correlations. This Perspective examines methodological foundations spanning structural causal models, disentangled causal representation learning, and techniques for interventional prediction and counterfactual reasoning on graphs. We analyse applications demonstrating clinical value across psychiatric diagnosis through brain network analysis, cancer subtyping via multi-omics causal integration, continuous physiological monitoring with mechanistic interpretation, and drug recommendation correcting prescription bias. These advances establish foundations for patient-specific Causal Digital Twins, enabling in silico clinical experimentation, with integration of large language models for hypothesis generation and causal graph neural networks for mechanistic validation. Substantial barriers remain, including computational requirements precluding real-time deployment, validation challenges demanding multi-modal evidence triangulation beyond cross-validation, and risks of causal-washing where methods employ causal terminology without rigorous evidentiary support. We propose tiered frameworks distinguishing causally-inspired architectures from causally-validated discoveries and identify critical research priorities, including hierarchical causal discovery exploiting biological modularity, integration of diverse data modalities, and regulatory pathways for systems making causal rather than purely associational claims. Success requires balancing theoretical ambition with empirical humility, computational sophistication with clinical interpretability, and transformative vision with uncompromising validation standards.
keywords:
causality, graphs, healthcare, interpretability, machine learning, time-series
1 Introduction
In 2019, a widely deployed risk prediction algorithm for the allocation of healthcare resources systematically underestimated the severity of the disease for Black patients by relying on healthcare costs as a proxy for health needs, resulting in discrimination affecting millions of patients annually [1]. That same year, a diabetic retinopathy screening system with specialist-level laboratory performance (>>90% sensitivity) faced critical failure when deployed in community clinics. The model struggled with environmental distribution shift, rejecting 21% of patient images as ungradable due to lighting and protocol differences rarely seen in training data, thereby creating workflow bottlenecks rather than improving care. These failures exemplify the triple crisis of AI in healthcare: models collapse under distribution shift, perpetuate discriminatory patterns embedded in historical data, and struggle to provide reliable mechanistic insight [2, 3]. The paradox of pneumonia risk prediction starkly illustrates this brittleness: models learnt to classify asthma patients as lower risk because they received aggressive early treatment in training hospitals, a spurious correlation that could lead to dangerous under-treatment when deployed elsewhere [4]. The question facing healthcare AI is no longer whether systems can achieve impressive accuracy in retrospective test sets, but whether they can survive clinical deployment across diverse institutions and maintain both effectiveness and equity.
This brittleness stems from a fundamental mismatch between what machine learning optimises and what healthcare requires. Clinical decisions inherently require understanding causal mechanisms, will this drug help this patient? Would this patient have developed complications under alternative treatment?, yet current models learn statistical associations that may reflect institutional practices, equipment characteristics, patient selection, or confounding variables rather than biological reality [5, 6]. The COVID-19 pandemic exposed these limitations with devastating clarity, where predictive models trained on historical data failed catastrophically when confronted with a novel pathogen and rapidly evolving clinical practices, highlighting the urgent need for methods that generalise based on causal invariance rather than statistical regularities [7, 8]. The consequences extend beyond accuracy metrics to patient harm, as models that encode spurious correlations can recommend interventions effective in training environments but ineffective or harmful in deployment settings. Moreover, mounting evidence demonstrates that AI systems trained on historical data perpetuate or amplify existing healthcare disparities, with purely statistical fairness metrics proving insufficient to address discrimination transmitted along causal pathways [9, 10, 11, 12].
Randomised controlled trials remain the gold standard for estimating causal effects, yet they address population-average treatment effects under specific protocols rather than patient-specific responses, cannot explore the combinatorial space of multi-drug interactions, and are infeasible for many clinically relevant questions due to ethical, temporal, or cost constraints. Causal graph methods extend this inferential capacity by enabling counterfactual reasoning for individual patients, leveraging observational data where randomisation is impossible, and scaling causal analysis to high-dimensional molecular and physiological networks beyond experimental reach.
The distinction between correlation and causation maps directly to Pearl’s Causal Hierarchy, which organises reasoning into three levels of increasing inferential power [13, 14]. Level 1 (Association) addresses observational queries through conditional probabilities $P(Y|X)$, answering “what is?” questions via pattern recognition, the domain where standard machine learning excels. Level 2 (Intervention) concerns the effects of actions, formalised using the do-operator $P(Y|\text{do}(X))$, addressing “what if we do?” questions essential for treatment planning. Level 3 (Counterfactual) queries alternative outcomes under hypothetical conditions, answering “what would have been?” questions critical for personalised medicine and retrospective analysis. Healthcare demands information at Levels 2 and 3, predicting treatment responses, comparing therapeutic strategies, and identifying causal disease mechanisms, yet standard models operate solely at Level 1, identifying statistical patterns without understanding the causal structure [15]. The Causal Hierarchy Theorem demonstrates that these levels form a strict hierarchy where information at higher levels cannot be derived from lower levels without additional causal assumptions [13]. This fundamental limitation explains why purely associational machine learning models, regardless of architectural sophistication or scale of training data, fail to answer interventional and counterfactual questions essential for robust clinical decision-making.
Biomedical systems naturally form networks, molecular interactions, brain connectivity, metabolic pathways, protein complexes, and disease comorbidity patterns, making graph representations the natural framework for encoding biological relationships [16, 17]. Graph neural networks extend traditional graph analysis beyond fixed topological measures by learning representations directly from graph-structured data through iterative message passing, where nodes aggregate information from neighbours via learnable neural transformations that capture both local motifs and global network properties [18, 19, 20]. This learning paradigm has demonstrated substantial advantages over handcrafted graph features in biomedical applications, from molecular property prediction on chemical graphs to psychiatric diagnosis from brain connectivity networks [21, 22, 23]. Yet standard GNNs inherit supervised learning’s fundamental limitation: they optimise predictive performance by exploiting any statistical pattern in training data, whether reflecting genuine biological mechanisms or spurious correlations from confounding, selection bias, or batch effects, treating all network edges equally regardless of underlying causal structure [24, 25].
The convergence of causal inference with graph neural networks addresses this challenge. Recent theoretical advances in causal representation learning provide a path forward.By explicitly modelling causal structures within graph architectures, we can develop models that identify genuine therapeutic targets, predict treatment responses under interventions never observed in training data, and discover disease mechanisms rather than spurious correlations [26, 27]. This convergence is particularly timely given the four concurrent developments. First, the accumulation of multi-modal biomedical data, from single-cell sequencing that captures 20,000+ gene expression profiles to continuous physiological monitoring that generates terabytes of temporal dynamics, provides unprecedented resolution for causal discovery. Datasets now capture interventional outcomes and temporal dependencies that were previously unavailable [28, 29, 30]. Second, regulatory frameworks increasingly demand explainable AI in healthcare, which requires not merely predictions, but a mechanistic understanding that clinicians can verify against domain expertise and biological knowledge [31, 32]. Third, documented failures of correlation-based models under distribution shift have created urgency for methods achieving robust generalisation through causal invariance. Fourth, the integration of causal inference with fairness constraints enables rigorous identification and mitigation of discrimination transmitted along specific causal pathways, providing tools for distinguishing legitimate predictive factors from discriminatory proxies, a distinction central to anti-discrimination law yet absent from traditional machine learning approaches [11, 33].
This Perspective examines causal graph neural networks, methods that integrate causal inference principles with graph-based deep learning architectures, as a methodological framework for transforming AI in healthcare from reactive pattern recognition to proactive causal medicine. Unlike previous papers that addressed GNNs in healthcare [34] or causal inference more broadly [14, 35] separately, we focus specifically on their integration, methods in which both graph structure and causal reasoning prove essential for clinical impact. We analyse how causal principles enhance GNN architectures for healthcare tasks, moving beyond correlation to mechanism, and examine current applications that demonstrate clinical utility across psychiatric diagnosis, cancer treatment, physiological monitoring, and drug discovery. We also identify fundamental challenges in scaling these methods to real-world deployment, including computational complexity, validation without ground truth, fairness under distribution shift, and regulatory considerations. Our analysis reveals both transformative promise and substantial barriers: while CIGNNs achieve superior generalisation across healthcare institutions, identify biologically plausible mechanisms, and maintain fairness properties under distribution shift, they face computational requirements precluding real-time application, difficulty validating causal claims from observational data, and gaps between methodological sophistication and clinical interpretability.
The result of this research is the development of patient-specific Causal Digital Twins, dynamic computational models built on causal GNN frameworks that enable clinicians to perform in silico experiments before clinical intervention. The digital twin paradigm has emerged as a transformative framework for personalised medicine, offering dynamic computational representations of individual patients that evolve through continuous data integration [36, 37]. Recent conceptualisation identifies four essential capabilities that distinguish medical digital twins from conventional simulation models: explainability through mechanistic rather than black-box architectures, intervenability enabling virtual testing of therapeutic scenarios, learnability allowing patient-specific model refinement over time, and diversability supporting uncertainty quantification through ensemble approaches [36]. Early clinical successes, notably the artificial pancreas for type 1 diabetes management, demonstrate feasibility within circumscribed physiological systems [38, 39]. However, extending digital twins to complex multisystem diseases requires methodological foundations that current data-driven approaches cannot provide. Causal graph neural networks could supply the missing theoretical infrastructure. Structural causal models enable the interventional reasoning essential for virtual treatment testing, counterfactual computation supports patient-specific outcome prediction, and invariant causal learning can potentially ensure robust generalisation across healthcare contexts. The integration of causal inference with graph-based patient modelling thus can allow us to transform digital twins from conceptual aspiration to methodologically grounded clinical tools. Imagine a clinician treating advanced cancer who could load the patient’s multi-omics profile, brain imaging, and clinical history into such a system, then simulate five drug combinations to predict effects on specific tumour pathways, toxicity risks, and progression-free survival, identifying optimal personalised therapy before administering a single dose. This vision, while ambitious, rests on methodological foundations now being established through causality-inspired architectures for interventional prediction, counterfactual generation, and robust multi-modal integration. Realising this potential requires not only technical innovation but fundamental rethinking of how we evaluate and deploy AI in medicine. We need to shift from predictive accuracy on retrospective test sets to causal validity under prospective deployment, from statistical fairness metrics to interventional equity guarantees, and from black-box pattern recognition to mechanistic interpretability verified against biological knowledge. Success requires coordinated advances across scalable causal discovery algorithms, validation frameworks suited to healthcare’s observational constraints, integration strategies for multi-modal multi-scale biological data, and regulatory pathways for systems making causal rather than purely associational claims. The path from proof-of-concept demonstrations to clinical implementation is neither straight nor simple, demanding collaboration between computational scientists, clinicians, biomedical researchers, and policymakers united by the recognition that healthcare AI’s next decade will be defined by which approaches achieve not just impressive metrics but genuine clinical trust through mechanistic understanding.
Causal graph neural networks address these fundamental limitations through principled causal inference frameworks (Fig. 1). By explicitly modelling causal structures within graph architectures, these methods identify invariant biological mechanisms rather than spurious correlations, enabling robust deployment across heterogeneous clinical settings and supporting mechanistic treatment optimisation through counterfactual reasoning. This Perspective proceeds in three stages. Section 2 establishes the theoretical foundations connecting structural causal models to graph neural network architectures. Section 3 presents methodological advances organised by clinical capability, disentanglement, intervention prediction, counterfactual generation, robustness, and fairness, providing the technical building blocks. Section 4 demonstrates clinical impact, showing how these methods translate to diagnosis, prognosis, treatment selection, and monitoring across healthcare domains. Section 5 synthesises these advances toward Causal Digital Twins, articulating how the methodological components combine to enable patient-specific in silico experimentation. Section 6 addresses the substantial barriers, computational, validation, and translational, that must be overcome for clinical deployment. We note that Sections 2–4 primarily synthesise existing methods and empirical evidence, while Section 5 articulates an aspirational framework, Causal Digital Twins, whose full realisation remains prospective. Section 6 addresses the substantial methodological, computational, and translational barriers separating current capabilities from this vision.
Figure 1: Clinical failure modes of traditional machine learning and solutions through causal graph neural networks. A, Clinical failure modes demonstrate fundamental limitations of correlation-based models in healthcare deployment. A.1, Vulnerability to distribution shift: a diabetic retinopathy screening system achieves 94% accuracy at the training institution (Hospital A) but performance collapses to 73% at deployment sites (Hospital B) due to spurious correlations with site-specific imaging protocols and patient demographics rather than causal disease mechanisms. Feature distributions show that spurious features (red, site-specific) undergo a distributional shift, while causal features (blue, disease markers) remain invariant across environments. A.2, Lack of mechanistic insight: traditional models operate as black boxes, learning associations between age, treatment, and comorbidity without identifying the underlying causal pathways, precluding clinical interpretability and mechanistic validation. Question marks indicate uncertain causal relationships that cannot be disentangled from observational data alone. A.3, Counterfactual blindness: Associational models trained on observational data (Treatment A, 87% five-year survival) cannot simulate alternative treatment scenarios (Treatments B and C) marked with prohibition symbols, as they lack the structural causal knowledge required for interventional reasoning, which is essential to personalised treatment selection. B, Causal GNN solutions address these limitations through principled causal inference. B.1, Causal invariance learning: multi-environment optimisation identifies stable causal relationships (blue nodes with shield protection) whilst suppressing environment-specific spurious correlations (red nodes, faded), ensuring robust generalisation under distribution shift through invariant risk minimisation. B.2, Causal digital twin for counterfactual generation: integration of multi-modal patient data (shown as concentric rings) constructs patient-specific structural causal models encoding disease mechanisms. Application of the do-operator enables simulation of unobserved treatment scenarios, generating counterfactual predictions (Treatment B: 73%, Treatment C: 95% five-year survival) through interventional inference rather than requiring empirical observations, facilitating treatment optimisation for individual patients.
2 The causal graph framework
2.1 Why biological systems demand graph-structured causal models
Biological systems exhibit causal relationships organised as networks across multiple scales of organisation. At the molecular level, gene regulatory networks encode transcriptional control in which regulatory proteins causally influence the expression of target genes, with modern single-cell RNA sequencing that captures the expression profiles of 20,000+ genes simultaneously [28]. Protein-protein interaction networks represent physical binding relationships that mediate cellular signalling, with causal perturbations propagating through these networks to generate phenotypic responses [16]. At the tissue level, brain networks comprise hundreds of regions with functional and structural connectivity encoding causal influences between neural populations, where activity in one region can drive or modulate activity in connected regions [40]. At the population level, disease comorbidity networks represent relationships where large-scale data mining of electronic health records and genetic association data are integrated into knowledge graphs to reveal how conditions are linked through shared molecular features [17]. This inherent graph structure makes neural network architectures that explicitly represent and reason about network topology the natural framework for biomedical machine learning.
The mathematical foundation connecting causal inference to graph representations is based on Structural Causal Models (SCMs), which support a rigorous formalisation of cause-effect relationships [13]. An SCM $\mathcal{M}:=\langle U,V,F,P(U)\rangle$ specifies exogenous variables $U$ determined outside the model, endogenous variables $V$ defined within the model, structural equations $F=\{f_{1},\ldots,f_{m}\}$ where each $V_{i}=f_{i}(\text{pa}(V_{i}),U_{i})$ with $\text{pa}(V_{i})$ denoting the parent set of $V_{i}$, comprising all variables that directly cause $V_{i}$ in the causal graph, and probability distribution $P(U)$ over exogenous variables [41]. This formulation naturally induces a directed acyclic graph $\mathcal{G}(\mathcal{M})$ where nodes represent variables and directed edges encode functional dependencies through the parent sets $\text{pa}(V_{i})$. The joint distribution factorises according to the Markov property: $P(V_{1},\ldots,V_{m})=\prod_{i=1}^{m}P(V_{i}|\text{pa}(V_{i}))$ [13]. Interventions in this framework correspond to “graph surgery,” modifying the causal graph by removing incoming edges to intervened variables and setting their values, fundamentally altering the generative process [13, 42]. This graph-theoretic representation of causation aligns precisely with the message-passing architecture of graph neural networks, enabling implementation of causal reasoning through differentiable neural operations.
A critical modelling choice underlying any causal graph concerns the semantic definition and resolution of nodes [43, 44]. In biomedical applications, nodes may represent entities at vastly different levels of abstraction such as individual genes versus pathways, single neurons versus brain regions, specific metabolites versus metabolic modules. This granularity is not merely a technical convenience but fundamentally determines which causal relationships can be identified and what interventions the graph implies [45]. Causal structure discovered at one resolution may not transfer to finer or coarser scales, as relationships that hold between brain regions might decompose into opposing sub-regional effects, while molecular interactions may aggregate into emergent pathway-level causation invisible at the gene level. Throughout this Perspective, we explicitly note the resolution at which different methods operate, as this choice carries direct consequences for both the biological interpretation of discovered causal structures and the experimental designs capable of validating them.
A related question concerns the epistemic status of existing biological network representations. Gene regulatory networks encode regulatory relationships that carry causal interpretation, perturbing transcription factor X should alter expression of target gene Y. However, GRNs inferred from observational expression data may conflate correlation with causation, with edges reflecting shared upstream regulators or cellular state confounding rather than direct regulation. Recent large-scale benchmarking on single-cell perturbation data reveals that even state-of-the-art inference methods struggle to recover true causal relationships when validated against interventional ground truth [46]. Systems biology models, comprising ordinary differential equations encoding metabolic and signalling dynamics, occupy different epistemic ground. These models encode mechanistic hypotheses derived from biochemical knowledge, enzyme kinetics, binding affinities, reaction stoichiometry, and thus represent causal mechanisms more explicitly. Yet parameters fitted to observational data may encode environment-specific rather than invariant relationships. Protein-protein interaction networks used by methods such as PDGrapher as ”causal graph approximations” [47] contain both direct causal interactions and indirect associations through protein complexes. We advocate treating biological networks as candidate causal structures requiring validation through interventional evidence, not established ground truth. The CiGNN framework exemplifies this by using causal discovery algorithms to identify genuinely causal physiological relationships rather than assuming database networks encode causation [48].
Standard GNNs, although naturally suited to graph-structured data, learn to exploit whatever statistical patterns maximise predictive performance regardless of causal validity or biological plausibility. In brain imaging, global arousal states create spurious correlations between all regions that GNNs readily exploit; in molecular networks, batch effects create artificial associations between unrelated proteins; in clinical data, treatment selection bias generates correlations between drug prescriptions and outcomes that reflect physician decision-making rather than therapeutic efficacy [24, 25]. These spurious patterns often provide a strong predictive signal in training environments, but fail catastrophically under distribution shift, precisely the brittleness observed in clinical deployment. Causality-inspired GNN architectures address this through architectural constraints, training objectives, and inductive biases that direct learning towards invariant causal relationships rather than environment-specific correlations. The DisC framework, for example, employs dual encoders to explicitly disentangle causal features that consistently predict outcomes across environments from spurious features that correlate with outcomes only in specific settings [49]. The CI-GNN architecture for brain networks uses conditional mutual information from a Granger causality perspective to identify connections representing genuine neural communication rather than shared global fluctuations [49, 50]. These approaches transform GNNs from being powerful function approximators that learn arbitrary predictive patterns to becoming mechanistic models that capture biologically grounded causal relationships.
2.2 From association to intervention to counterfactuals
Standard supervised learning excels at associational queries, given covariates $X$, predict the outcome $Y$ by learning $P(Y|X)$ from observational data. This suffices for many pattern recognition tasks: identifying malignant tumours from images, predicting patient deterioration from vital signs, or classifying disease subtypes from gene expression. However, clinical decision-making fundamentally requires interventional reasoning: what happens if we administer treatment $T$? This requires estimating $P(Y|\text{do}(T))$, which differs from $P(Y|T)$ whenever confounding exists, and confounding pervades observational healthcare data through indication bias, selection effects, and unmeasured factors [14, 13].
Graph-structured causal models enable interventional prediction through Pearl’s do-calculus, which provides three rules for transforming interventional distributions into estimable observational quantities under specified graph conditions (Box 1) [13]. Rule 1 allows ignoring observations on variables that don’t affect the outcome through paths independent of the intervention. Rule 2 permits converting interventions to observations when appropriate conditional independence holds. Rule 3 enables removing interventions when they do not affect the outcome through certain graph paths. The interventional Variational Graph Auto-Encoder (iVGAE) implements these operations through differentiable neural layers, incorporating interventional transformations in both encoder and decoder networks to ensure that the learnt representations encode causal rather than associational relationships [42]. When the model identifies a connection between two proteins or two brain regions, it learns whether this connection would persist under interventions, whether it represents a causal pathway or a spurious correlation.
Counterfactual reasoning operates at the highest level of Pearl’s hierarchy, addressing patient-specific questions: would this patient have developed complications if we had chosen alternative treatment? Reasoning of this kind is essential for precision medicine, treatment optimisation, and retrospective analysis of clinical decisions [13]. The calculation of counterfactuals requires a three-step process, including abduction (inferring unobserved variables, including patient-specific factors, from observed data), action (modifying the causal model to reflect alternative intervention), and prediction (computing outcomes under the modified model) [41]. The CLEAR framework implements counterfactual generation for graphs through variational autoencoders that learn distributions over causally plausible graph modifications. This could allow, for example, the exploration of alternative biological configurations with some mechanistic validity [51]. In cancer treatment planning, this makes it possible to simulate how different therapeutic targets would affect tumour evolution, identifying vulnerabilities that might not be apparent from observational data alone.
Throughout this Perspective, we distinguish methods by the strength of their causal claims: causally-inspired architectures that employ causal concepts without explicit causal claims, causally-structured methods that estimate causal quantities under stated assumptions, and causally-validated approaches whose claims are corroborated through multi-modal evidence (see Section 6.3 for formal definitions).
Patient-specific Causal Digital Twins, dynamic computational models where clinicians perform in silico experiments to test interventions before clinical administration could integrate the multi-omics profile of a patient (genomics, transcriptomics, proteomics), longitudinal imaging that captures organ structure and function, clinical history including treatments and responses, and knowledge graphs encoding established biological mechanisms [28, 29]. The underlying causal GNN architecture would learn patient-specific parameterisations of general biological mechanisms, enabling simulation of therapeutic interventions through the do-operator framework by setting treatment variables, propagating effects through learnt causal pathways, and predicting outcomes at molecular, cellular, and phenotypic levels. For an oncology patient, the digital twin might simulate five drug combinations, predicting effects on specific oncogenic pathways, normal tissue toxicity, and progression-free survival, thereby identifying optimal personalised therapy before administering a single dose. For a neuropsychiatric patient, it could simulate neuromodulation protocols, predicting which brain circuits would be affected and how symptoms would respond. This vision, while ambitious, rests on methodological foundations now being established of GNN architectures for interventional prediction, counterfactual generation frameworks for exploring alternative scenarios, invariant learning methods ensuring robust generalisation, and multi-modal integration strategies combining diverse data modalities through shared causal structure [26, 52]. Figure 2A illustrates these theoretical distinctions, depicting the conceptual framework through which traditional graph neural networks may achieve comparable associational performance but are expected to exhibit substantial degradation for interventional and counterfactual queries essential for clinical decision-making.
Box 1: Formal foundations of causal graph neural networks Structural Causal Models An SCM $\mathcal{M}:=\langle U,V,F,P(U)\rangle$ defines the data-generating process, where exogenous variables $U=\{U_{1},\ldots,U_{n}\}$ represent unmodelled background factors, endogenous variables $V=\{V_{1},\ldots,V_{m}\}$ are defined by structural equations $V_{i}=f_{i}(\text{pa}(V_{i}),U_{i})$, and $P(U)$ governs exogenous randomness [13]. The induced causal graph $\mathcal{G}(\mathcal{M})$ contains directed edges $V_{j}\to V_{i}$ for each $V_{j}\in\text{pa}(V_{i})$. Interventions and the do-operator Intervention $\text{do}(X=x)$ modifies the SCM by replacing the structural equation for $X$ with $X:=x$, removing incoming edges in the graph. The post-intervention distribution $P(Y|\text{do}(X=x))$ differs from observational $P(Y|X=x)$ whenever confounding exists [13]. Graph neural network implementation GNNs implement message-passing: $\mathbf{h}_{i}^{(l+1)}=\sigma\left(\mathbf{W}^{(l)}\mathbf{h}_{i}^{(l)}+\sum_{j\in\mathcal{N}(i)}\mathbf{M}^{(l)}\mathbf{h}_{j}^{(l)}\right)$, where $\mathbf{h}_{i}^{(l)}$ is the representation of node $i$ at layer $l$, $\mathcal{N}(i)$ are neighbours, and $\mathbf{W},\mathbf{M}$ are learnable parameters. This architecture naturally aligns with causal graphs when edges represent causal relationships and message-passing implements causal influence propagation [42]. Causal identifiability $P(Y|\text{do}(X))$ is identifiable if it can be expressed in terms of the observational distribution $P(V)$ given the causal graph $\mathcal{G}$. The do-calculus provides three rules for such expressions [13]: $P(Y|\text{do}(X),Z,W)$ $=P(Y|\text{do}(X),W)\text{ if }Y\perp_{\mathcal{G}_{\overline{X}}}Z|X,W$ (1) $P(Y|\text{do}(X),\text{do}(Z),W)$ $=P(Y|\text{do}(X),Z,W)\text{ if }Y\perp_{\mathcal{G}_{\overline{X}\underline{Z}}}Z|X,W$ (2) $P(Y|\text{do}(X),\text{do}(Z),W)$ $=P(Y|\text{do}(X),W)\text{ if }Y\perp_{\mathcal{G}_{\overline{X}\overline{Z(W)}}}Z|X,W$ (3) where $\mathcal{G}_{\overline{X}}$ denotes the graph with edges into $X$ removed, $\mathcal{G}_{\underline{Z}}$ has edges out of $Z$ removed, and $\perp$ indicates d-separation. Rule 1 permits ignoring observations on variables that do not affect the outcome through paths independent of the intervention; Rule 2 allows converting interventions to observations when appropriate conditional independence holds; Rule 3 enables removing interventions when they do not influence the outcome through certain graph paths. Counterfactual computation For unit $u$ and intervention $X=x$, the counterfactual outcome $Y_{x}(u)$ is computed via: (1) Abduction: infer $U=u$ from observed data; (2) Action: modify SCM to $\mathcal{M}_{x}$ with $X:=x$; (3) Prediction: evaluate $Y$ under $\mathcal{M}_{x}$ with $U=u$ [41]. This enables patient-specific predictions unavailable from population-level interventional distributions.
Figure 2: Conceptual framework illustrating theoretical advantages of causal over associational inference across Pearl’s causal hierarchy. A, Theoretical dissociation between associational and interventional performance predicted by causal hierarchy theory. The “causal gap” quantifies expected performance loss when models trained on observational data are applied to interventional queries, reflecting the fundamental limitation that associational patterns may not persist under intervention. B, Idealised robustness to simulated unmeasured confounding. Conceptual curves illustrate the theoretical expectation that traditional methods degrade more steeply as unmeasured confounder strength increases, while causal architectures that explicitly model confounding structure are expected to maintain robustness. Confounding strength parameterised as maximum risk ratio for unmeasured confounder associations with treatment and outcome following sensitivity analysis frameworks [53]. Shaded regions represent idealised uncertainty bands illustrating expected variance patterns. Causally-Inspired GNN (Tier 1) denotes architectures employing causal concepts such as invariance regularisation to improve robustness without explicit causal claims. Causally-Validated GNN (Tier 3) denotes methods whose causal claims are corroborated through multi-modal evidence triangulation. See Section 6.3 for formal tier definitions.
3 Methodological advances
The translation of causal theory into practical healthcare applications requires architectures that address specific clinical capabilities. We organise methodological advances by the clinical capabilities they enable, disentangling genuine causal mechanisms from confounding, predicting interventional effects without experimental data, generating patient-specific counterfactuals, achieving robust generalisation across heterogeneous settings, and ensuring fairness through causal rather than purely statistical criteria, discussing within each capability the technical approaches that address it. From a complementary technical perspective, methods reviewed here rest on distinct theoretical foundations that determine their causal assumptions and evidentiary requirements. Intervention-invariance methods (Sections 3.1, 3.4) assume that causal features produce stable predictions across environments while spurious correlations shift with context; these require multi-environment data but not necessarily explicit causal graphs. Structural causal methods (Sections 3.2, 3.3) implement do-calculus through neural architectures, requiring explicit graph specification, causal sufficiency, and faithfulness assumptions; these provide formal identification guarantees but demand stronger prior knowledge. Knowledge-guided methods incorporate biomedical priors, protein-protein interactions, gene regulatory networks, anatomical atlases, as structural constraints without formal causal identification; these leverage domain expertise but inherit the epistemic status of their knowledge sources, which may conflate correlation with causation (Section 2.1). Table 2 maps individual methods to their specific assumptions within this taxonomy.
3.1 Disentangling causal signal from confounding
Biomedical datasets are rife with spurious correlations that mislead standard machine learning models. Standard GNNs, although naturally suited to graph-structured data, learn to exploit whatever statistical patterns maximise predictive performance regardless of causal validity or biological plausibility. In brain imaging, global arousal states create spurious correlations between all regions that GNNs readily exploit [54]; in molecular networks, batch effects create artificial associations between unrelated proteins [55]; in clinical data, treatment selection bias generates correlations between drug prescriptions and outcomes that reflect physician decision-making rather than therapeutic efficacy [24, 25]. The challenge is to decompose observed graphs into causal and spurious components, a problem addressed through information-theoretic disentanglement that maximises mutual information between causal features and outcomes and minimises spurious correlations [56].
The DisC (Debiasing via Disentangled Causal Substructure) framework operationalises this separation through a dual-encoder architecture [49]. Rather than treating graphs monolithically, DisC maintains parallel GNN encoders: one trained to extract causal features that consistently predict outcomes across different environments, another capturing environment-specific spurious patterns. The contrastive learning objective enforces the independence between these representations, ensuring that causal and spurious features remain disentangled. For psychiatric diagnosis, this architectural separation was proven to be immensely effective. Traditional GNNs identified widespread hyperconnectivity in the cerebellar regions of depressed patients, a finding inconsistent with the clinical understanding that depression primarily influences affective rather than motor circuits.
The CI-GNN (Causality-Inspired Graph Neural Network) advances this concept through Granger causality-inspired conditional mutual information, providing principled measures of causal strength for graph edges [50]. The model identifies influential subgraphs that demonstrate genuine causal connections through disentangled subgraph-level representation learning. CI-GNN’s causal decomposition correctly identified sparse cerebellar involvement while highlighting strong connections around the left rectus, a region previously identified as critical for antidepressant treatment response prediction [50]. Evaluated on the REST-meta-MDD dataset (n=1,604), this alignment with established neurobiology demonstrates that causal disentanglement provides mechanistically meaningful insights, not merely improved prediction. In autism spectrum disorder evaluations on the ABIDE dataset, CI-GNN revealed tight connections within the somatomotor network consistent with motor impairments characteristic of ASD, findings replicated across independent cohorts and suggesting genuine causal relationships rather than dataset-specific artefacts [50]. The clinical imperative was that clinicians needed to know which connections to target through neurostimulation, pharmacotherapy, or behavioural interventions, not merely that connectivity is abnormal. Although developed for other applications, IGCL-GNN extends these principles through multi-objective optimisation that simultaneously enhances causal features and suppresses environment-specific spurious patterns, potentially being transferrable to help with robust clinical deployment which requires balancing predictive performance on current datasets with generalisation to unseen environments [56].
These findings depend on the chosen parcellation resolution. CI-GNN operates on 116-node brain atlases where each node aggregates thousands of neurons; causal relationships at this mesoscale resolution reflect functional circuit organisation rather than individual neuronal connectivity, with direct implications for which neurostimulation or pharmacological interventions could target the identified causal pathways. Figure 3 illustrates these architectural principles applied to functional neuroimaging data. While the DisC and iVGAE frameworks were originally developed on synthetic benchmark datasets, their conceptual translation here to brain connectivity analysis could demonstrate how causal disentanglement and interventional prediction could address the specific confounding challenges in clinical neuroimaging, separating genuine neural circuit dysfunction from global arousal states, scanner artefacts, or site-specific acquisition protocols.
Figure 3: Canonical causal graph neural network architectures illustrated for conceptual brain connectivity analysis. While the underlying methods were developed on synthetic benchmark datasets, we illustrate their application to functional neuroimaging to demonstrate translation potential for clinical contexts. A, The DisC (Debiasing via Disentangled Causal Substructure) framework [49] achieves causal disentanglement through a dual-encoder architecture. An edge mask generator partitions input graphs into causal and spurious subgraphs, with separate GNN modules encoding each into disentangled representations. The loss comprises two components: $\mathcal{L}_{CE}$, a weighted cross-entropy loss that up-weights samples where the bias GNN struggles, directing the causal encoder toward genuinely predictive features; and $\mathcal{L}_{GCE}$, a generalised cross-entropy loss that amplifies gradients on easily-classified samples, directing the bias encoder toward environment-specific shortcuts. Causal features ($Z_{C.1}$$Z_{C.4}$) produce stable predictions across environments while spurious features ($Z_{S.1}$$Z_{S.3}$) are explicitly identified and discarded. B, The iVGAE (interventional Variational Graph Auto-Encoder) framework [42] implements Pearl’s do-operator through differentiable neural layers. The interventional encoder maps observed graphs to a causal latent space $Z$, while the interventional decoder incorporates do-layers that simulate graph surgery. The output $P(\mathbf{D}|\textit{do}(\mathbf{X}))$ represents the interventional distribution of downstream variables $\mathbf{D}$ given intervention $\textit{do}(\mathbf{X}=x)$, enabling prediction of effects that persist under manipulation rather than spurious correlations reflecting confounding. Applied to brain networks, such architectures could distinguish genuine neural communication pathways from global arousal-induced correlations, identifying intervention targets for neuromodulation or pharmacotherapy.
3.2 Predicting interventional effects without experimental data
Precision medicine requires not just prediction, but mechanistic explanation, identifying the specific biological features driving a diagnosis. The RC-Explainer (Reinforced Causal Explainer) uses reinforcement learning to discover which edges of a graph to retain to maximise the causal information flow [57]. Rather than computing interventions symbolically, it learns a policy for selecting explanatory subgraphs, with each candidate evaluated not by correlation but by causal impact. In molecular property prediction, this approach identified causal functional groups (e.g., toxicophores) determining mutagenicity, distinguishing active substructures from inert scaffolds, mechanistic insights missed by correlation-based methods that cannot distinguish causal features from spurious environmental associations [57].
For blood pressure monitoring, the CiGNN framework demonstrates the clinical utility of interventional prediction by first using the neural-accelerated causal discovery (FCI algorithm) to identify physiological signals causally related to blood pressure variation, then constructing spatiotemporal GNN architectures guided by the discovered causal structure [48, 58]. This causal feature selection reduces dimensionality from 222 initial wearable features to a sparse set of direct and indirect causal indicators while improving both accuracy and interpretability, identifying, for example, that specific PPG morphological features (e.g., pulse arrival time, waveform characteristics) causally influence blood pressure through arterial stiffness mechanisms.
The iVGAE (interventional Variational Graph Auto-Encoder) provides theoretical grounding for interventional GNNs by incorporating interventional layers in both encoder and decoder networks, ensuring that learnt representations explicitly encode causal rather than merely associational relationships [42]. When the model identifies a connection between proteins or brain regions, it learns whether this connection represents a causal pathway that would persist under intervention rather than a spurious correlation arising from confounding. For drug discovery, PDGrapher leverages protein-protein interaction networks as approximations of causal graphs, operating under causal sufficiency assumptions to predict combinatorial therapeutic targets [47]. Evaluated on LINCS L1000 chemical perturbation data across nine cell lines (A549, MCF7, PC3, HT29, A375, and others), PDGrapher outperformed competing methods by 4.5–13.4% in detecting ground-truth targets, demonstrating that even imperfect causal structures, PPI networks contain both direct causal effects and indirect associations, provide superior guidance compared to purely correlational feature selection.
3.3 Generating counterfactuals for personalised medicine
Counterfactual reasoning addresses the highest level of Pearl’s causal hierarchy: what would have happened to this patient under alternative treatment? Such patient-specific predictions are essential for precision medicine, treatment optimisation, and retrospective analysis of clinical decisions [13]. The CLEAR (Counterfactual Learning on Graphs) framework approaches counterfactual generation through variational autoencoders that learn distributions over possible graph modifications [51]. Rather than generating arbitrary alternative graphs, CLEAR learns which modifications are causally plausible, which edges could change, maintaining validity relative to the data distribution. In molecular discovery, this leads to exploration of how structural modifications would affect molecular properties, identifying functional groups that drive toxicity or efficacy, vulnerabilities invisible in observational data alone [51].
A challenge in counterfactual reasoning is that interventions can create graph configurations far outside the training distribution. When simulating disruption of a critical hub protein, we can query a configuration likely never appearing naturally, and thus unrepresented in training data. Traditional neural networks can fail potentially under such a distribution shift. For epileptic seizure prediction, dynamic graph models like DIGLN capture the temporal evolution of brain networks [59], future extensions could enable in silico testing of interventions, simulating how disrupting specific neural circuits would affect seizure propagation patterns, guiding neurosurgical planning or neurostimulation protocols.
The CXGNN (Causal Explainer based on GNN-Neural Causal Models) framework provides a rigorous quantification of counterfactual effects through neural implementations of structural causal models [60]. By parameterising each structural equation as a neural network, CXGNN learns arbitrary non-linear causal relationships and stays true to the formal guarantees of SCM-based counterfactual computation. For mild cognitive impairment progression, CXGNN identified that hypertension, arrhythmia, and sex emerged as critical causal factors, with notable sex-specific differences: female patients showed stronger causal links between vascular factors and cognitive decline, while male patients exhibited more pronounced metabolic effects [60, 61]. These mechanistic insights enable personalised risk assessment and targeted early intervention strategies. These counterfactual generation methods form the computational engine of Causal Digital Twins, the ability to simulate patient-specific responses to interventions never administered in reality constitutes the core functionality enabling in silico clinical experimentation.
3.4 Achieving robustness across heterogeneous clinical settings
While disentanglement methods (Section 3.1) achieve robustness by explicitly separating causal from spurious representations, invariant learning approaches pursue the same objective through a complementary mechanism: optimisation constraints that enforce feature stability across environments without requiring explicit architectural separation. Disentanglement provides interpretable causal-spurious decomposition and invariant learning provides deployment guarantees with potentially simpler architectures. Some methods, such as IGCL-GNN, integrate both principles.
Healthcare AI deployment requires models that generalise across institutions with different patient demographics, clinical protocols, and measurement technologies, precisely the distribution shift conditions where correlation-based models fail catastrophically. The fundamental principle underlying invariant learning is that causal relationships remain stable across contexts while spurious correlations change (Figure 2B) [62]. In medical terms, the biological mechanisms by which drugs treat diseases should be consistent across populations, even though confounding factors such as demographics, comorbidities, and healthcare practices vary widely. This principle provides a powerful criterion for distinguishing causal from spurious patterns that hold across diverse environments and are likely causal.
The IGCL-GNN framework relies on invariant causal learning through multi-objective optimisation that enhances causal features and suppresses environment-specific spurious patterns [56]. Though developed for industrial sensor networks, it provides a template for robust clinical deployment requires balancing competing objectives such as maximising predictive performance on current datasets and also ensuring generalisation to unseen environments. In multi-site brain imaging studies, CI-GNN achieved mean accuracy of 0.71–0.73 across 17 acquisition sites compared to 0.64–0.69 for baseline methods, with consistent improvements in 13 of 17 sites despite individual site accuracy ranging from 0.60 to 0.85, demonstrating the value of learning causal invariances rather than site-specific correlations [49, 50]. This stability under distribution shift is not just a technical achievement, but a prerequisite for clinical trust, models that collapse when moved between hospitals cannot be deployed safely.
The CRec (Conditional Causal Representation Learner) framework addresses the robustness of drug recommendation through disentanglement of therapeutic effects from prescription patterns [63]. For example, proton pump inhibitors might appear beneficial for cardiovascular outcomes in observational data, not due to a therapeutic effect but due to a confounding pattern that varies across institutions, namely co-prescription with aspirin for cardioprotection. CRec separates drug molecular representations into causal therapeutic components and spurious prescription patterns through interventional disentanglement. In an analysis of 47,326 patients with multiple chronic conditions, CRec identified mechanistically synergistic drug combinations that traditional methods missed because they were rarely co-prescribed, revealing genuine therapeutic opportunities obscured by prescribing traditions [63]. The DIR-GNN framework extends invariant learning through environment-based partitioning, explicitly learning which features remain predictive across all environments (causal invariants) versus environment-specific features (spurious correlations), then building predictions based solely on invariant features [64].
3.5 Ensuring causal fairness in clinical algorithms
Traditional fairness metrics based on statistical parity, equalising predictions across demographic groups, operate at Pearl’s Level 1 (association), yet clinical deployment decisions inherently require Level 2 (intervention) reasoning [13, 11]. The critical question for healthcare systems is not whether predictions correlate equally across demographics but whether predictions remain stable under hypothetical interventions on sensitive attributes, a distinction with profound implications for equity. Statistical fairness may equalise predictions across groups while preserving discriminatory causal mechanisms embedded in the data-generating process.
Interventional fairness requires that the predicted outcome distributions remain invariant under interventions on sensitive attributes: $P(\hat{Y}=y|\text{do}(A=a))=P(\hat{Y}=y|\text{do}(A=a^{\prime}))$ for all outcomes $y$ and sensitive attribute values $a,a^{\prime}$ [9, 10]. This criterion extends naturally to path-specific fairness, where effects along specific causal pathways are constrained. For the Standard Fairness Model with sensitive attribute $A$, confounders $Z$, mediators $W$, and outcome $Y$, we decompose the total effects into Natural Direct Effect (NDE), Natural Indirect Effect (NIE), and Natural Spurious Effect (NSE), representing effects transmitted through different causal pathways [11]. A causally $S$-fair predictor for pathway set $S$ ensures that the pathways in $S$ transmit no discriminatory signal and preserves the causal structure for pathways not in $S$, preventing undesirable bias amplification [65].
The RaVSNet framework demonstrates the clinical necessity of causal bias correction by explicitly modelling the discrepancy between historical prescription habits and therapeutic efficacy [66]. Using Bayesian network causal discovery (GIES algorithm), RaVSNet quantifies how institutional protocols and co-occurrence patterns influence medication choices independently of patient-specific benefit. The system generates causal effect matrices to adjust recommendation probabilities, filtering out spurious associations driven by data availability rather than medical necessity. On MIMIC-III and MIMIC-IV, RaVSNet achieved Jaccard scores of 0.573 and 0.497 respectively (3–7% improvement), while its causal inference module specifically mitigates the structural confounding in electronic health records, an important step for preventing the propagation of historical inequities that often manifest as data sparsity or noise in underserved populations [66].
Path-specific decomposition enables explicit deliberation about equity costs and benefits [33]. Analysis reveals that removing direct effects of sensitive attributes often incurs minimal accuracy loss while substantially reducing disparity, suggesting that certain forms of discrimination can be eliminated at low cost. Conversely, constraining indirect effects through legitimate mediators (education, occupation) imposes higher accuracy costs, reflecting genuine predictive value. This transparency enables stakeholders such as regulatory bodies, healthcare institutions, and patient advocacy groups, to make informed decisions about acceptable fairness-utility trade-offs based on quantitative causal evidence rather than speculation. For clinical deployment, interventional fairness addresses a critical limitation: models fair by statistical criteria may exhibit discrimination when deployed across different healthcare contexts experiencing a distribution shift. Causally-constrained predictors maintain fairness properties by learning invariant causal relationships rather than site-specific correlations [49, 50].
4 Clinical impact across domains
The translation of causal graph methods into clinical practice represents a convergence of methodological innovation with pressing healthcare needs. These applications span from individual patient diagnosis to population-level disease surveillance, demonstrating how causal inference transforms raw biomedical data into actionable clinical insights. Success depends not only on technical sophistication, but also on the addressing of fundamental clinical questions about disease mechanisms, treatment responses, and intervention strategies.
4.1 Diagnosis: from correlation to mechanism
4.1.1 Mechanistic brain network analysis
Brain network analysis exemplifies the critical need for causal over correlational approaches. The human brain comprises approximately 86 billion neurons that form trillions of connections, creating one of the most complex networks in nature. Traditional graph-theoretical methods, while foundational, often fall short in capturing the high-dimensional and dynamic nature of brain connectivity [67]. The brain naturally forms a graph where regions of interest serve as nodes and functional connectivity defines edges, yet correlation-based analysis frequently identifies spurious connections driven by global brain states, arousal levels, or measurement artefacts rather than genuine neural communication pathways [50].
Causality-Inspired Graph Neural Networks tackle this through an interpretable architecture that fundamentally reimagines brain network analysis [49, 50]. Rather than treating all connectivity equally, CI-GNN learns disentangled latent representations where only the causal factor induces the explanatory subgraph related to diagnosis. This methodology leverages conditional mutual information from a Granger causality perspective to measure causal influence, providing a principled distinction between genuine neural dysfunction and epiphenomenal changes. For major depressive disorder, CI-GNN identified strong connections around the left rectus, a region critical for the response to antidepressant treatment, while showing sparser connections in the cerebellum [50]. This finding contradicts spurious cerebellar hyperconnectivity found by association-based methods, aligning instead with clinical understanding that depression primarily involves affective and cognitive functions rather than motor function. Mechanistic insights such as these prove crucial for developing targeted interventions through neurostimulation, pharmacotherapy, or behavioural therapy.
The BPI-GNN framework advances psychiatric diagnosis through prototype learning that identifies biologically meaningful subtypes of psychiatric disorders [68]. Beyond discriminating patients from healthy controls, the model revealed subtypes with distinct clinical profiles. In major depressive disorder (REST-meta-MDD, n=1,604), identified subtypes exhibited distinct patterns on HAMD-17 clinical subscales including suicide risk, psychomotor retardation, and somatic symptoms, suggesting that network-based biomarkers could inform clinical characterisation. For schizophrenia (SRPBS, n=184), subtypes corresponded to differences in PANSS symptom profiles across positive, negative, and general psychopathology domains [68]. The model achieved superior performance against 11 popular brain network classification methods across three psychiatric datasets (REST-meta-MDD, ABIDE with n=1,064, and SRPBS), demonstrating both technical excellence and clinical relevance.
For epileptic seizure prediction, the Dynamic Instance-level Graph Learning Network (DIGLN) captures patient-specific causal relationships between intracranial electroencephalogram channels that evolve over time [59]. By learning both the organisation of epileptic networks and their temporal evolution, DIGLN demonstrated the ability to evaluate multiple overlapping patterns of brain functional change, particularly in connectivity between the visual and sensorimotor domains [69]. In 5-fold cross-validation on the Freiburg iEEG dataset (9 patients, 45 seizures), this patient-specific approach achieved 85.7% recall and 91.7% precision (F1: 87.0%), demonstrating the potential of dynamic graph learning for seizure prediction, though external validation on larger cohorts remains necessary.
4.1.2 Multi-omics cancer subtyping
Cancer represents a medical challenge where the integration of multiple data modalities is not only beneficial, but essential to understand the mechanisms of the disease. Single-omic approaches may lack the precision required to establish robust associations between molecular-level changes and phenotypic traits [70]. Tumours are characterised by genomic instability, epigenetic dysregulation, transcriptomic reprogramming, and proteomic alterations, with each layer providing complementary information about cancer biology.
Causal methodology transforms cancer classification by focusing on causally related genes rather than on all differentially expressed features. The CautionGCN framework employs a causal multi-head autoencoder for causal feature selection, isolating characteristic genes with genuine causal relationships to cancer subtypes [71]. This approach addresses a fundamental problem in cancer genomics, where thousands of genes show differential expression, but only a subset causally drives tumorigenesis. By constructing causal gene-gene graphs that capture driver relationships under assumptions of causal sufficiency and faithfulness (Table 2).
The gene-level resolution of MoCaGCN’s causal graph carries important validation implications where each identified causal driver can, in principle, be experimentally tested through CRISPR knockout or pharmacological inhibition of the specific gene product. This contrasts with pathway-level causal analyses where nodes aggregate multiple genes, creating ambiguity about which molecular targets should be perturbed to validate the discovered causal relationship.
The identified causal genes provide more than classification accuracy. The MOGONET framework takes advantage of both omics features and correlations among samples described by similarity networks, using the Correlation Discovery Network to explore cross-omics correlations in the label space [72]. This approach recognises that different omics layers may capture distinct aspects of cancer biology, with genomics revealing mutational drivers, transcriptomics showing pathway activation, and proteomics indicating functional consequences. By learning these multi-level representations jointly, MOGONET achieved superior performance in breast invasive carcinoma subtype classification while identifying layer-specific biomarkers.
The clinical value extends beyond improved diagnostic accuracy to mechanistically informed treatment planning. Understanding which molecular alterations causally drive disease progression, rather than merely correlating with it, enables the rational design of targeted therapies and the identification of vulnerabilities that are amenable to pharmacological intervention. The shift from correlation-based feature selection to causal driver identification represents a fundamental advance in precision oncology.
4.2 Prognosis: predicting what interventions prevent
The Knowledge-aware Multi-label Network (KAMLN) addresses a significant limitation in postoperative complication prediction by explicitly modelling dependencies between complications rather than treating them as independent events [73]. The framework utilizes a heterogeneous knowledge graph depicting clinical relationships such as ”lung infection causes pleural effusion,” which in turn is linked to atelectasis [73]. In a cohort of 593 lung cancer surgery patients, KAMLN achieved a micro-AUC of 0.664, demonstrating that integrating prior medical knowledge can stabilize predictions for rare complications—such as pulmonary embolism and deep vein thrombosis—where traditional single-label models often fail [73]. The model’s use of SHAP analysis and a five-layer graph neural network architecture provides a transparent approach to understanding the clinical features that drive multi-label risk, identifying factors like operation duration and the number of dissected lymph nodes as critical predictors across multiple complication groups [73]. While further external validation is required before direct clinical application, the explicit modelling of these causal chains facilitates more accurate and fine-grained risk assessment for surgical patients [73].
For the progression of mild cognitive impairment to dementia, the CXGNN framework, which employs generative causal explainers based on the GNN-Neural Causal Model methodology, moves beyond static snapshots to identify the primary causal factors driving the transition [60]. Analysis revealed that hypertension, arrhythmia, and sex emerged as critical variables, with notable sex-specific differences in the pathways of cognitive decline. Female patients showed stronger causal links between vascular factors and cognitive decline, while male patients exhibited more pronounced effects of metabolic factors. This mechanistic understanding enables a personalised risk assessment that goes beyond simple risk scores, providing actionable insights for an early intervention tailored to individual risk profiles. Clinicians can target specific causal pathways rather than applying uniform interventions, potentially improving efficacy and reducing unnecessary treatment burden.
4.3 Treatment selection: mechanism-based therapy
Drug recommendation in the era of polypharmacy requires understanding not just individual drug effects but also complex interactions and patient-specific factors. The Conditional Causal Representation Learner tries to tackle spurious correlations that plague traditional recommendation systems [63]. For example, proton pump inhibitors might appear beneficial for cardiovascular outcomes not due to their therapeutic effect but because they are prescribed to patients who receive aspirin for cardioprotection. CRec disentangles these confounded relationships by separating drug molecular representations into causal therapeutic components and spurious prescription patterns. Through causal intervention to eliminate spurious effects, recommendations become genuinely mechanism-focused. In a study of 47,326 patients with multiple chronic conditions, CRec identified drug combinations that traditional methods failed to identify because they were rarely co-prescribed despite mechanistic synergy [63]. One notable discovery involved the combination of a diabetes medication with an Alzheimer’s drug based on shared metabolic pathways, a combination now in clinical trials showing promising results for both conditions.
RaVSNet improves the precision of the recommendation by explicitly modelling and correcting data distribution bias through Causal Effect Inference tasks [66]. Using Bayesian network causal discovery (GIES algorithm), RaVSNet quantifies how clinical factors (diagnoses, symptoms) influence medication choices independent of simple co-occurrence patterns. The system generates causal effect matrices that adjust recommendation probabilities to reflect genuine therapeutic associations rather than spurious statistical correlations. On MIMIC-III and MIMIC-IV datasets, RaVSNet achieved Jaccard scores of 0.573 and 0.497 respectively, improvements over state-of-the-art baselines, while the causal framework provides a principled mechanism for identifying and correcting confounding biases in retrospective electronic health records [66].
For drug discovery, PDGrapher uses protein-protein interaction networks as approximations of causal graphs, operating under the assumption that there are no unobserved confounders to predict combinatorial therapeutic targets [47]. This causally-inspired approach outperforms traditional methods in identifying effective perturbagens across nine cell lines with chemical perturbations. The identified causal targets provide more than simple efficacy predictions; by solving the inverse problem of perturbagen discovery, the model reveals combinatorial dependencies where targeting multiple specific proteins is required to shift a unique disease state to a healthy profile. These mechanistic insights could allow the rational design of combination therapies that exploit biological structure rather than trial-and-error. Consequently, the clinical workflow transforms from ”drug prescribed for patients like you” to ”drug targeting your causal mechanisms,” representing a fundamental shift toward mechanism-based precision medicine.
4.4 Continuous monitoring: real-time causal inference
Biomedical monitoring generates vast streams of temporal data where the distinction of causal signals from noise is essential for clinical decision-making. The challenge becomes particularly acute in intensive care settings where hundreds of variables are monitored continuously, creating high-dimensional time series with complex interdependencies. Traditional monitoring systems generate frequent false alarms, with studies showing that up to 92% of ICU alarms are clinically irrelevant, leading to alarm fatigue and potentially missed critical events [74].
The CiGNN framework for cuffless blood pressure estimation exemplifies how causal methods transform physiological monitoring [48, 58]. Rather than using all available features from electrocardiogram and photoplethysmogram signals, CiGNN first applies causal discovery using FCI and CGNN algorithms to identify features causally related to variation in blood pressure. This causal feature selection reduces dimensionality from more than 200 potential features to 23 causal indicators, dramatically improving both computational efficiency and physiological interpretability. The causal relationships discovered provide physiological insights beyond prediction. CiGNN identified that specific PPG morphological features causally influence blood pressure through arterial stiffness mechanisms, while certain heart rate variability patterns affect blood pressure through autonomic regulation. These mechanistic insights enable personalised monitoring strategies in which patients with autonomic dysfunction require different monitoring parameters than those with primarily vascular pathology.
In high-stakes clinical monitoring, missing data is often the reality which spatio-temporal modelling has to overcome. To this end, the Casper framework can provide a proof-of-concept to addresses the challenge of unknown confounders that create non-causal shortcut backdoor paths [75]. While evaluated on sensor networks, the framework’s use of frontdoor adjustment offers a methodology to eliminate spurious correlations, such as those created by hospital environmental factors influencing both monitoring frequency and patient physiological parameters, while preserving genuine causal relationships [75]. The Spatiotemporal Causal Attention module of the framework enforces sparse, interpretable causality among embeddings, with theoretical guarantees of convergence to gradient-based explanations [75]. By demonstrating substantial improvements in imputation quality and reduced variance in complex multi-site datasets, Casper provides a robust foundation for more accurate clinical data reconstruction [75]. This enhanced accuracy potentially improves the prediction of patient deterioration and reduces the likelihood of unnecessary interventions triggered by imputation errors, effectively translating causal discovery into improved clinical outcomes.
The REASON framework extends causal analysis to system-level monitoring, crucial to understanding cascading failures in complex medical systems [76]. By modelling both intra-level and inter-level non-linear causal relationships, REASON can trace how local perturbations propagate through interconnected physiological systems. In sepsis monitoring, REASON identified how initial inflammatory responses cascade through the cardiovascular, respiratory, and renal systems, enabling early intervention before multi-organ failure develops. This system-level causal analysis represents a paradigm shift from monitoring individual parameters to understanding how perturbations propagate through coupled biological networks, enabling proactive rather than reactive clinical management.
Table 1: Comparative analysis of representative causal graph neural network methods by clinical capability
Task Methods Key Innovation Clinical Impact
Diagnostic subtyping CI-GNN [50], BPI-GNN [68] Disentangled causal subgraph identification via conditional mutual information Psychiatric subtypes with distinct treatment responses; accuracy 0.71–0.73 with cross-site stability across 17 institutions
Cancer classification MoCaGCN[77], CautionGCN[71] Causal gene graph construction integrating prior biological knowledge 89 causal genes vs. 1,000+ correlational features; identified KRAS-metabolic axis for combination therapy
Seizure prediction DIGLN [59] Patient-specific dynamic causal connectivity learning 85.7% recall, 91.7% precision on Freiburg iEEG (n=9); mechanistic understanding of propagation patterns
Complication prediction KAMLN [73] Heterogeneous knowledge graph encoding clinical dependencies between complications Micro-AUC 0.664 for multi-label prediction; SHAP analysis identifies lymph node dissection as a significant risk factor
Drug recommendation CRec [63], RaVSNet [66] Disentanglement of therapeutic effects from prescription bias Identified mechanistically synergistic combinations (CRec); 4–7% accuracy improvement over baselines with reduced DDI rates (RaVSNet)
Combination therapy PDGrapher [47] Inverse causal inference on protein-protein interaction networks Direct prediction of combinatorial perturbagens for phenotype reversal; 13.3% improvement in identifying therapeutic targets from interventional data
Blood pressure monitoring CiGNN [48] FCI-based causal feature selection for wearable sensors 200+ features to 23 causal indicators; mechanistic interpretation of cardiovascular regulation
ICU data imputation Casper [75] Frontdoor adjustment to block unknown spatiotemporal confounders Significantly reduced MAE/MSE; methodological template for handling non-causal shortcuts in continuous monitoring
Cognitive decline CXGNN [60] Neural causal models for counterfactual reasoning Sex-specific causal pathways; personalized risk assessment for MCI-to-dementia progression
  • Abbreviations: CI-GNN, Causality-Inspired Graph Neural Network; BPI-GNN, Brain Prototype-Inspired GNN; MoCaGCN, Multi-omics Causal Graph Convolutional Network; DIGLN, Dynamic Instance-level Graph Learning Network; KAMLN, Knowledge-Aware Multi-label Network; CRec, Conditional Causal Representation Learner; RaVSNet, Relevance-Aware Visit Similarity Network; PDGrapher, Perturbation-Driven Grapher; CiGNN, Causality-informed Graph Neural Network; Casper, Causality-Aware Spatiotemporal Imputation; CXGNN, Causal Explainer GNN; FCI, Fast Causal Inference; MCI, Mild Cognitive Impairment; AUC, Area Under Curve.
Table 2: Causal assumptions and evidentiary status of methodological classes
Method Class Representative Methods Required Assumptions Causal Claim Tier
Disentanglement DisC, CI-GNN, IGCL-GNN Environment diversity sufficient for separation; causal features invariant across environments Separation of stable vs. unstable features Tier 1
Interventional prediction iVGAE, RC-Explainer, PDGrapher Causal sufficiency (no unobserved confounders); faithfulness; correct graph specification $P(Y|\text{do}(X))$ estimation Tier 2
Counterfactual generation CLEAR, CXGNN, CSA Causal sufficiency; faithfulness; correct functional form of structural equations Patient-specific $Y_{x}(u)$ computation Tier 2–3
Causal discovery CiGNN (FCI), MoCaGCN Faithfulness; causal Markov condition; acyclicity (or known cyclic structure) Graph structure learning Tier 2–3
Fairness constraints RaVSNet, path-specific methods Correct causal graph; no unmeasured confounding of sensitive attribute pathways Path-specific effect removal Tier 2
  • Tier 1: causally-inspired (architectural use of causal concepts without explicit causal claims); Tier 2: causally-structured (causal estimates validated against external knowledge); Tier 3: causally-validated (novel causal discoveries with multi-modal triangulation). See Section 6.3 for tier definitions. Causal sufficiency assumes all common causes of measured variables are observed. Faithfulness assumes statistical independencies reflect only graphical d-separation.
5 Towards causal digital twins
Table 3: Mapping methodological capabilities to clinical applications and digital twin components
Capability (§3) Clinical Application (§4) Digital Twin Role (§5) Key Methods
Causal disentanglement Psychiatric subtyping; cancer driver identification Isolating patient-specific mechanisms DisC, CI-GNN, IGCL-GNN
Intervention prediction Drug combination discovery; treatment selection Virtual therapeutic testing iVGAE, RC-Explainer, PDGrapher
Counterfactual generation Personalised prognosis; retrospective analysis Patient-specific outcome simulation CLEAR, CXGNN, CSA
Invariant learning Cross-institutional deployment; monitoring robustness Generalisation across clinical contexts CaT-GNN, DIR-GNN
Causal fairness Bias-corrected recommendations; equitable risk prediction Equity-preserving simulation RaVSNet, path-specific constraints
Table 3 summarises how the methodological capabilities developed in Section 3 and validated across clinical domains in Section 4 converge toward the integrative vision of patient-specific causal simulation. Digital twins have transitioned from manufacturing origins to healthcare applications, representing a paradigm shift toward dynamic, patient-specific computational models [36, 78]. Unlike static predictive models, digital twins maintain bidirectional connections with their physical counterparts, continuously updating representations as new patient data become available. The conceptual framework articulated by Emmert-Streib and colleagues identifies four capabilities including explainability through mechanistic model structures interpretable in biological terms, intervenability enabling simulation of therapeutic interventions before clinical administration, learnability through iterative refinement as patient trajectories unfold, and diversability supporting probabilistic exploration of alternative health states [36]. Applications spanning cardiovascular medicine [79], immunology [37], and oncology [80] have demonstrated domain-specific utility, while closed-loop systems such as the artificial pancreas illustrate mature clinical deployment for circumscribed physiological control [38].
Yet a critical methodological gap constrains broader translation. Digital twins are not inherently causal models, current implementations predominantly rely on either physics-based simulations of specific organs or data-driven models that learn patient-specific parameters without formal causal grounding. Physics-based approaches achieve mechanistic interpretability but scale poorly to multisystem diseases involving thousands of interacting molecular and physiological variables. Data-driven approaches scale efficiently but inherit the fundamental limitation of associational machine learning in which they cannot distinguish causal mechanisms from spurious correlations, precluding reliable intervention simulation when deployed beyond training distributions. Causal graph neural networks resolve this tension by combining the scalability of neural architectures with the theoretical guarantees of structural causal models. The do-calculus provides formal machinery for interventional prediction, counterfactual computation enables patient-specific outcome simulation, and invariant causal learning ensures robustness across heterogeneous clinical contexts. This methodological synthesis establishes the foundation for Causal Digital Twins, patient-specific computational models in which clinicians perform rigorous in silico experiments grounded in causal rather than merely associational inference.
5.1 Patient-specific mechanistic simulation
Causal inference and graph learning can together enable a transformative vision for precision medicine. Patient-specific Causal Digital Twins represent dynamic computational models in which clinicians perform in silico experiments to test interventions before clinical administration. Such systems would integrate a patient’s multi-omics profile encompassing genomics, transcriptomics, proteomics, and metabolomics; longitudinal imaging capturing organ structure and function over time; clinical history including treatments, responses, and adverse events; and knowledge graphs encoding established biological mechanisms from decades of biomedical research [28, 29, 17]. The underlying causal GNN architecture would learn patient-specific parameterisations of general biological mechanisms, enabling simulation of therapeutic interventions through the do-operator framework. By setting treatment variables, propagating effects through learnt causal pathways, and predicting outcomes at molecular, cellular, and phenotypic levels, the system generates mechanistically grounded predictions unavailable from purely associational models.
Consider a clinician treating advanced non-small cell lung cancer with EGFR mutations. The patient’s Causal Digital Twin would encode the specific mutational landscape from whole-genome sequencing, transcriptomic profiles that reveal pathway dysregulation, proteomic data showing aberrant signalling cascades, and imaging demonstrating tumour burden and metastatic sites. The clinician would, under this framework, be able to simulate five therapeutic strategies through the digital twin. First-generation EGFR inhibitors targeting the primary mutation, combination therapy pairing EGFR inhibition with immune checkpoint blockade, metabolic targeting based on observed Warburg effect signatures, anti-angiogenic therapy addressing vascular supply, or sequential therapy alternating between mechanisms to prevent resistance. For each simulated strategy, the digital twin generates predictions conditional on the assumed causal structure, estimating effects on specific oncogenic pathways using learnt causal relationships from the patient’s molecular data, estimates toxicity risks by propagating drug effects through normal tissue networks, projects progression-free survival by modelling tumour evolution under selection pressure, and identifies optimal personalised therapy before administering a single dose. This capability transforms treatment planning from empirical trial-and-error to mechanistically informed optimisation.
Current methodological advances establish foundations for realising this vision. Interventional GNNs such as iVGAE enable simulation of treatments never observed in training data by implementing do-calculus through differentiable neural operations [42]. Counterfactual generators, including CLEAR and CXGNN, explore alternative scenarios while maintaining biological plausibility, computing patient-specific predictions through the abduction-action-prediction framework [51, 60]. Multi-modal integration frameworks such as MOGONET and MoCaGCN combine diverse data modalities through shared causal structure, learning representations that capture biological mechanisms spanning molecular to phenotypic scales [72, 77]. Fairness constraints ensure equitable simulation by identifying and mitigating discriminatory causal pathways, preventing digital twins from perpetuating biases embedded in training data [11].
5.2 The LLM-Causal GNN synergy
The integration of large language models with causal GNNs creates a powerful hybrid architecture addressing complementary aspects of clinical reasoning. Large language models excel at processing unstructured clinical data, including free-text notes, radiology reports, pathology descriptions, and biomedical literature, extracting entities, relationships, and temporal patterns from narrative text [29, 81]. However, LLMs lack rigorous causal inference capabilities and cannot distinguish genuine causal relationships from spurious correlations in their training data. Causal GNNs provide the complementary capability of validating and quantifying causal hypotheses using structured multi-omics data, medical imaging, and biological knowledge graphs, but require well-defined graph structures and struggle with unstructured information.
The synergistic workflow proceeds through hypothesis generation followed by mechanistic validation. An LLM continuously monitors a patient’s electronic health record, parsing clinical notes, laboratory results, medication orders, and vital signs. When the system detects a prescribed drug that coincides temporally with emerging adverse symptoms, the LLM proposes a causal hypothesis linking drug exposure to observed effects. This hypothesis, expressed as a potential causal pathway, is then submitted to a Causal GNN that accesses protein-protein interaction networks capturing drug-target relationships and off-target effects, pharmacovigilance databases containing population-scale adverse event reports, and the patient’s genomic data revealing drug metabolism variants. The Causal GNN assesses whether the proposed relationship satisfies causal identifiability criteria, computes interventional distributions predicting effects under different dosing strategies, performs sensitivity analysis quantifying robustness to unmeasured confounding, and returns a mechanistically validated assessment in seconds [13, 14]. If the relationship proves causal, the system alerts the clinician with a mechanistic explanation of the pathway linking the drug to adverse effect, a quantitative estimate of effect magnitude based on the patient’s molecular profile, and alternative medication recommendations avoiding the problematic mechanism. This hybrid approach combines the clinical reasoning capabilities of LLMs, including natural language understanding and medical knowledge retrieval, with the mechanistic validation of Causal GNNs, providing rigorous quantification of causal effects and biological plausibility checking. The synergy proves particularly powerful for pharmacovigilance, where novel adverse drug reactions often first appear in clinical narratives before accumulating sufficient structured reports for traditional signal detection. By enabling real-time causal evaluation of hypotheses extracted from unstructured text, the LLM-Causal GNN architecture could potentially identify safety signals years earlier than current passive surveillance systems [82].
Such hybrid architectures must contend with well-documented LLM limitations in biomedical contexts. Systematic evaluations reveal substantial hallucination rates when LLMs generate scientific references, ranging from 29% for GPT-4 to over 90% for some models, with poor adherence to specified inclusion criteria even for non-hallucinated outputs [83]. Comprehensive benchmarking across biomedical natural language processing tasks demonstrates that while LLMs excel at reasoning-intensive question answering, they underperform fine-tuned domain-specific models in information extraction, with qualitative analyses revealing prevalent missing information and inconsistent outputs [84]. These findings underscore that causal GNNs contribute not merely complementary functionality but essential grounding which by encoding structured biomedical knowledge and enforcing mechanistic constraints, can anchor LLM-generated hypotheses to validated biological relationships, potentially mitigating the hallucination risks that currently limit clinical deployment of pure LLM approaches [85].
5.3 Implementation requirements
Translating the Causal Digital Twin vision to clinical reality requires coordinated advances across computational infrastructure, data integration, validation frameworks, and regulatory pathways. Computational requirements pose immediate challenges. Current causal GNN methods incur substantial computational overhead from causal discovery and multi-environment optimisation. While CI-GNN achieves efficient inference (approximately 33ms per instance on 116-node brain networks using the AAL atlas), training requires grid search over hyperparameters across multiple independent runs [50], and realistic molecular networks contain 20,000+ genes and protein-protein interaction networks span hundreds of thousands of edges. Real-time clinical decision support demands inference times measured in seconds, not hours. Algorithmic innovations, including hierarchical causal discovery that exploits biological modularity, amortised inference approaches that train once but infer rapidly, and optimised neuromorphic hardware for graph operations, represent promising directions [76]. Cloud-based computational infrastructure with distributed graph processing capabilities could enable the necessary scale while maintaining clinical response times.
Multi-modal data integration pipelines must harmonise heterogeneous data sources spanning genomic sequences, gene expression matrices, proteomic mass spectrometry, metabolomic profiles, medical images in DICOM format, clinical notes in natural language, structured electronic health records, and continuous physiological monitoring streams. Each modality requires specialised preprocessing, quality control, and normalisation prior to integration through a shared causal structure [28, 29]. Temporal alignment proves particularly challenging when different modalities are collected at varying frequencies, from genome sequences measured once to physiological signals sampled thousands of times per second. Federated learning approaches that preserve privacy enable model training across institutions without centralising sensitive patient data, addressing both technical and regulatory constraints [52].
Clinical validation protocols must establish evidentiary standards for digital twin predictions. The multi-modal triangulation framework described earlier provides a starting point, combining biological plausibility assessment, replication across patient cohorts, natural experiments where available, and prospective validation studies comparing digital twin predictions to observed outcomes [13]. Regulatory frameworks for in silico experimentation remain underdeveloped. The FDA and EMA have established pathways for in silico clinical trials in the development of medical devices, but extending these to patient-specific treatment optimisation requires new guidance that addresses unique challenges of personalised mechanistic modelling [86]. Questions of liability when digital twin predictions prove inaccurate, intellectual property for patient-specific models, and equitable access to computationally intensive precision medicine tools demand policy attention. Success requires sustained collaboration between computational scientists developing causal inference methods, clinicians defining clinical requirements and validation criteria, biomedical researchers providing mechanistic knowledge, and regulatory bodies establishing appropriate oversight frameworks. The technical foundations are established; organisational and policy innovations will determine whether Causal Digital Twins transform from an aspirational vision to a clinical reality.
6 Challenges and path forward
6.1 Computational reality
Algorithmic innovations targeting biological structure offer promising solutions. Hierarchical causal discovery methods exploit the modular organisation of biological systems, first identifying causal relationships between subsystems such as organ networks or cellular pathways, then refining the structure within the module [76]. This approach dramatically reduces computational complexity by leveraging natural sparsity patterns where most biological components interact locally within functional modules rather than exhibiting all-to-all connectivity. The REASON framework demonstrates feasibility by modelling both intra-level and inter-level causal relationships, enabling system-level analysis of cascading failures in complex physiological networks [76]. Amortised inference represents another transformative direction, where models are trained once on diverse datasets and then perform rapid inference on new patients. Rather than discovering the de novo causal structure for each patient, the system learns to predict the causal relationships from the observed data through end-to-end differentiable architectures [60]. This enables real-time causal inference, which is feasible for clinical decision support. For applications where distribution shift is minimal and intervention prediction unnecessary, correlation-based approaches remain appropriate and more practical; causal methods are warranted specifically when the failure modes they address, brittleness under deployment shift, confounded treatment effects, counterfactual personalisation, are clinically relevant.
Hardware acceleration through neuromorphic computing and graph-specific processors offers complementary advances. Graph neural network operations exhibit inherent parallelism amenable to specialised hardware, with emerging architectures achieving orders-of-magnitude speedups for message-passing computations [42]. Mixing algorithmic innovation, and exploiting biological structure and amortised learning, with hardware acceleration through specialised computing substrates, establishes a viable path toward clinically deployable causal GNN systems operating at the scale and speed demanded by precision medicine applications.
[…]
6.2 Circumstances of failure
Recent large-scale empirical work directly challenges optimistic assumptions about causal generalisation. Nastl and Hardt evaluated 16 prediction tasks across healthcare, employment, education, and social policy domains, each with multiple deployment environments permitting rigorous out-of-distribution testing [87]. Without exception, predictors using all available features, regardless of causality, achieved superior in-domain and out-of-domain accuracy compared with predictors restricted to causal features. Even the absolute accuracy drop from training to deployment domains was no better for causal predictors than for models exploiting all correlational patterns. Causal machine learning methods specifically designed for domain generalisation did not perform better than standard predictors trained on causal features alone, while causal discovery algorithms either failed to execute on real-world datasets or selected variables offering no advantage over expert-curated selections. These findings suggest that causal approaches provide benefit primarily when non-causal features experience substantial concept shift between domains; when distributional changes are modest, the information loss from discarding potentially predictive features outweighs robustness gains.
Sample size constraints pose particular challenges for causal GNNs in neuroimaging applications. Brain connectivity studies typically involve high-dimensional graphs comprising hundreds to thousands of nodes with relatively small patient cohorts, leading to a higher risk of overfitting [67]. Neural network models exhibit larger variance than conventional machine learning methods on small datasets, and regularisation strategies effective for standard GNNs may not transfer to causal architectures with additional structural constraints. The BrainGB benchmark revealed that sophisticated GNN architectures can exhaust memory on large-scale datasets while providing minimal advantages over shallow baseline models on small datasets where overfitting dominates [88]. The fundamental tension between model complexity required for causal reasoning and data availability in clinical neuroimaging remains unresolved.
Unobserved confounding constitutes a limitation that architectural innovation cannot overcome (Figure 2B). Causal conclusions from observational data rest on the assumption of no unmeasured confounding, yet this assumption is empirically untestable and rarely satisfied in healthcare settings [15]. When treatment decisions depend on unmeasured clinical judgement, patient preferences, or institutional practices, even sophisticated causal GNN architectures may learn spurious relationships that appear causally valid within their structural assumptions but reflect residual confounding. The CausalBench benchmark, evaluating causal discovery methods against interventional single-cell perturbation data with known ground truth, demonstrated that state-of-the-art methods struggle to recover true regulatory relationships [46]. Substantial performance gaps persist between low-intervention and high-intervention experimental settings, suggesting that methods validated on observational benchmarks may not transfer to genuine interventional prediction.
These empirical limitations underscore that causal GNNs should be deployed with appropriate epistemic humility. The methods excel when specific failure modes, distribution shift between training and deployment, confounded treatment effects requiring debiasing, or counterfactual personalisation demanding patient-specific simulation, are clinically relevant and sufficiently severe to justify computational overhead. When these conditions are absent, correlation-based approaches remain appropriate and more practical.
6.3 The validation crisis
Establishing trustworthy causal conclusions from observational healthcare data constitutes perhaps the most fundamental challenge facing the field. Cross-validation and held-out test sets effectively evaluate predictive performance, but provide insufficient evidence to validate causal claims [89, 13]. The core problem is that standard machine learning evaluation metrics, including accuracy, area under the curve, and calibration measure association rather than causation. A model can achieve excellent predictive performance while encoding spurious correlations that would fail under intervention [90, 62]. Ground truth causal structures remain unknown for most biomedical systems, precluding direct validation of learnt causal graphs against known mechanisms.
Table 2 lays out the assumptions underlying each methodological class reviewed in this Perspective. Critically, violation of these assumptions, particularly causal sufficiency in the presence of unmeasured confounding, invalidates the causal interpretation of model outputs regardless of predictive performance. The tiered evidentiary framework provides a mechanism for calibrating confidence: Tier 1 methods make no causal claims requiring these assumptions; Tier 2 methods require assumption plausibility supported by domain knowledge; Tier 3 claims demand empirical validation that assumptions hold in the specific application context.
The validation crisis (Figure 2A) illustrates that methods clustered in the high-correlation, low-intervention quadrant achieve impressive retrospective metrics while failing the prospective deployment. This ”causal gap” quantifies the expected performance loss when correlation-based models are applied to interventional queries, validating the need for validation frameworks extending beyond cross-validation to natural experiments, propensity-matched analyses, and prospective intervention studies [89, 91].
Multi-modal evidence triangulation provides a rigorous framework for establishing causal validity through convergent evidence across complementary modalities [92, 89]. Biological plausibility assessment evaluates whether identified causal relationships align with established pathophysiological mechanisms and biochemical pathways, providing mechanistic coherence testing against decades of accumulated biomedical knowledge [93]. The MoCaGCN framework demonstrates this approach by integrating prior knowledge from the GeneMANIA database to constrain learnt causal gene graphs, ensuring that identified relationships respect established gene-gene interactions while remaining flexible to discover novel causal mechanisms [77]. Replication among independent cohorts tests whether causal relationships remain stable across diverse populations, healthcare systems, and temporal periods. The CiGNN blood pressure monitoring framework was evaluated across three datasets, an internal cohort (n=62), VitalDB surgical patients (n=205), and a maneuver dataset (n=38), achieving MAE of 3.2–4.2 mmHg for systolic and 1.9–2.8 mmHg for diastolic blood pressure, demonstrating consistent performance across healthy and clinical populations [48].
Validation design must also address the resolution at which causal claims are made. Under the interventionist framework, causal relationships are defined relative to interventions that could, in principle, be performed [45]. When computational nodes aggregate multiple biological entities like ”prefrontal cortex” comprising dozens of sub-regions, ”inflammation” encompassing multiple cytokine cascades, or ”metabolic reprogramming” spanning hundreds of metabolites, experimental validation becomes ambiguous. Interventions targeting specific sub-components may not test the aggregate causal claim, while interventions at the aggregate level may be infeasible or imprecise. This resolution mismatch between computational causal discovery and experimental validation represents a systematic challenge: causal graphs operating at pathway resolution require experimental designs that modulate entire pathways coherently, whereas gene-level graphs permit targeted perturbation experiments. Explicit specification of node semantics in causal graph reporting, analogous to intervention definitions in clinical trial protocols, would clarify what experiments could validate the claims and at what biological scale the causal relationships are asserted to hold.
Natural experiments and quasi-experimental designs provide powerful causal validation by exploiting situations where treatment assignment approximates randomisation through exogenous factors [94, 91]. Policy discontinuities, instrumental variables, and regression discontinuity designs enable the estimation of causal effects with minimal confounding, serving as benchmarks against which observational causal models can be validated. Prospective intervention studies, when feasible, constitute the gold standard by testing whether the predicted intervention outcomes materialise under actual treatment [89]. Sensitivity analyses quantify robustness to violations of causal assumptions, particularly unmeasured confounding, through E-value calculations and model specification uncertainty assessment [53, 95].
We propose tiered evidentiary standards that distinguish the strength of causal evidence. Tier 1 encompasses causally-inspired methods that employ causal concepts in architectural design but make no explicit causal claims about learnt relationships, such as invariance-based regularisation promoting features stable across environments [96]. Tier 2 includes causally-structured methods that learn causal structures or estimate causal effects validated through consistency with external knowledge rather than novel discovery, exemplified by cancer subtyping models recovering known oncogenic pathways [77]. Tier 3 represents causally-validated methods that claim discovery of previously unknown causal relationships, requiring comprehensive triangulation across biological plausibility with the proposed mechanism, replication across two or more independent cohorts, validation of natural experiments or prospective intervention studies, sensitivity analysis demonstrating robustness with E-values exceeding 2.0 for strong claims, and aggregate evidence scores exceeding 0.6 using quantitative integration frameworks [92].
6.4 From methods to clinical practice
The gap between methodological sophistication and clinical interpretability poses critical barriers to deployment. Clinicians require not only predictions but also mechanistic explanations that they can understand, verify against domain expertise, and integrate with clinical reasoning [32]. A significant challenge in the implementation of GNNs for biomarker discovery involves ensuring the interpretability and clinical relevance of identified patterns [67]. Computational representations learnt by neural networks often lack direct correspondence to clinical concepts, necessitating translation frameworks that bridge this semantic gap.
Integration of domain knowledge into model design represents a promising direction. Rather than learning causal structures purely from data, hybrid approaches incorporate established biological mechanisms as structural priors, constraining model architectures according to known pathway relationships [77]. The MoCaGCN framework demonstrates this by integrating gene-gene interaction networks from curated databases as graph topology priors, ensuring that learnt causal relationships respect established biological constraints while remaining flexible enough to discover novel mechanisms. Explainability frameworks must translate learnt causal representations into clinically meaningful concepts. The CI-GNN architecture achieves this by identifying sparse subgraphs corresponding to specific disease mechanisms, where highlighted regions and connections of the brain are directly linked to established neural circuits known to clinicians [49, 50].
Clinician-in-the-loop validation provides essential reality checking where domain experts evaluate whether computationally identified causal mechanisms align with clinical understanding and biological plausibility. This iterative refinement process, combining data-driven discovery with expert knowledge, proves critical for building clinical trust. Regulatory pathways for causal claims from observational data remain underdeveloped. The FDA has established frameworks for predictive algorithms but lacks specific guidance for systems making interventional or counterfactual claims based on causal inference from non-randomised data [86]. Developing appropriate oversight requires balancing innovation incentives with patient safety, establishing evidentiary standards sufficient to support causal claims without demanding infeasible validation burdens.
The risk of ”causal-washing”, where methods employ causal terminology without meeting rigorous evidentiary standards for causal claims, threatens field credibility. The appeal of causal framing is understandable given superior theoretical properties and alignment with clinical reasoning, yet the gap between causal aspiration and causal validation remains substantial across much of the literature. Several patterns characterise causal-washing in current practice. Terminological overreach occurs when methods use causal vocabulary, including ”causal representation,” ”causal discovery,” or ”causal reasoning,” while relying fundamentally on associational quantification without estimating interventional distributions or testing identifiability conditions [13, 62]. Insufficient validation arises when causal claims rest solely on predictive performance metrics without triangulating evidence from biological plausibility, replication, natural experiments, or sensitivity analyses. Assumption elision involves implementing causal methods while omitting explicit statements of the requisite assumptions, including faithfulness, causal sufficiency, and positivity, preventing the assessment of validity [26, 62].
Community-level practices can mitigate causal-washing through standardised reporting guidelines analogous to CONSORT for randomised trials or STROBE for observational studies [97]. Such guidelines should require an explicit statement of all causal assumptions with directed acyclic graphs when applicable, a specification of which causal quantities are being estimated, including the average treatment effect or conditional average treatment effect, reporting of sensitivity analyses quantifying robustness to assumption violations, and a clear demarcation between causal claims versus associational findings [14]. The transition from causal-washing to rigorous causal inference requires cultural change alongside methodological standards, where funding agencies prioritise validation studies over novel method proliferation, peer review demands explicit justification for causal terminology and rejects unsupported claims, and graduate training emphasises distinctions between causal inference, prediction, and description.
Box 3: Example of multi-modal validation framework The amyloid-$\beta$ hypothesis in Alzheimer’s disease Consider a causal GNN identifying a directed edge from amyloid-$\beta$ deposition to cognitive decline in Alzheimer’s disease, a relationship with validation complexities. Evidence triangulation: 1. Biological plausibility ($S_{\text{bio}}=+1$): Strong mechanistic support through the amyloid cascade hypothesis and decades of foundational neuroscience research. 2. Replication ($S_{\text{rep}}=+1$): Association replicates consistently across major cohorts, including ADNI, AIBL, and European EADC datasets with stable effect sizes. 3. Natural experiments ($S_{\text{quasi}}=-1$): Mendelian randomization studies using genetic variants affecting amyloid production demonstrate no causal effect on cognitive decline. 4. Prospective validation ($S_{\text{prosp}}=-1$): Multiple Phase III clinical trials of amyloid-reducing therapies show minimal cognitive benefit despite substantial plaque reduction. 5. Sensitivity analysis ($S_{\text{sens}}=0$): Moderate robustness to unmeasured confounding (E-value = 1.8) but substantial uncertainty regarding alternative causal structures. Aggregate evidence score: $\text{AES}=0.20(+1)+0.25(+1)+0.30(-1)+0.15(-1)+0.10(0)=0.00$ (4) Interpretation: Despite strong observational associations and biological plausibility, the null aggregate score could reflect contradictory evidence from high-quality natural experiments and prospective trials [92, 98]. This suggests the observational causal model may be confounded, with amyloid deposition potentially representing a biomarker of disease progression rather than a causal driver, or causality involving more complex dynamics, including threshold effects or critical developmental periods that the learned graph structure fails to capture. Clinical implication: Predictions based on this causal edge should be treated with appropriate caution, and amyloid reduction alone may not constitute an effective therapeutic strategy. This exemplifies how systematic evidence triangulation prevents overconfident causal claims despite compelling observational evidence, ultimately protecting patients from ineffective interventions whilst guiding research toward more productive therapeutic targets.
[…]
We note that healthcare presents additional methodological challenges where causal reasoning proves essential but which exceed the scope of this Perspective. Small-sample learning remains problematic for causal GNNs, as deep learning models require large datasets to avoid overfitting while neuroimaging studies typically involve high-dimensional graphs with limited patient cohorts [100]. Multi-centre generalisation, though addressed by invariant learning (Section 3.4), involves domain adaptation complexities requiring specialised harmonisation techniques beyond causal invariance alone [101]. Time-varying confounding in longitudinal treatment regimes demands methods such as marginal structural models and g-computation that extend beyond the cross-sectional causal graphs emphasised here [102]. Missing data mechanisms, ubiquitous in electronic health records, require causal models for principled imputation yet remain inadequately addressed in current causal GNN frameworks [103]. These directions warrant dedicated methodological attention as the field matures.
7 Conclusion
The integration of causal inference with graph neural networks addresses a fundamental limitation in healthcare artificial intelligence. The documented triple crisis of distribution shift that causes performance collapse across institutions, perpetuation of discriminatory patterns embedded in historical data, and lack of mechanistic interpretability that prevents clinical verification demand a paradigm change from correlation-based prediction to mechanism-based understanding [1, 2, 104]. This review has demonstrated that causality-inspired architectures, when rigorously implemented and validated, achieve superior generalisation across healthcare settings, identify biologically plausible mechanisms aligned with established pathophysiology, and maintain fairness properties under distribution shift through learning of invariant causal relationships rather than site-specific correlations [49, 50, 11].
Applications spanning psychiatric subtyping, cancer molecular characterisation, physiological monitoring, and drug discovery validate this approach across diverse clinical domains. The CI-GNN framework maintained stable diagnostic performance across institutions where traditional methods varied dramatically [50]. Multi-omics causal integration identified 89 driver genes from thousands of correlational features, accelerating precision oncology [77]. Causal feature selection transformed continuous monitoring from high-dimensional noise to mechanistically interpretable signals [48]. The methodological foundations for patient-specific Causal Digital Twins now exist, enabling in silico clinical experimentation before therapeutic intervention, with synergistic integration of large language models for hypothesis generation and causal graph neural networks for mechanistic validation, positioning the field at the forefront of multimodal artificial intelligence [29, 82].
However, substantial barriers constrain translation to clinical practice. Computational complexity precludes real-time deployment, training that requires hours on specialised hardware for moderately sized networks, while realistic molecular graphs span tens of thousands of nodes [49, 28]. Validation must evolve beyond predictive accuracy to establish causal claims through multi-modal evidence triangulation combining biological plausibility, cross-cohort replication, natural experiments, and sensitivity analyses [89, 92]. The field faces a growing risk of causal-washing, where methods employ causal terminology without meeting rigorous evidentiary standards, threatening scientific credibility and clinical trust. Tiered frameworks distinguishing causally-inspired architectures from causally-structured models and causally-validated discoveries prove essential for maintaining appropriate epistemic humility while enabling continued innovation.
Healthcare artificial intelligence is at an inflexion point. The question is no longer whether machine learning can achieve impressive metrics in retrospective datasets but whether systems can survive deployment across diverse clinical contexts while maintaining effectiveness, equity, and interpretability. Causal graph neural networks represent the only architectural paradigm that simultaneously addresses robustness through invariant learning, fairness through path-specific interventional constraints, and mechanistic understanding through explicit causal structure [26, 62, 10]. Success requires coordinated action on multiple fronts. Computational infrastructure investments must enable hierarchical causal discovery exploiting biological modularity, amortised inference achieving clinical response times, and neuromorphic hardware acceleration for graph operations at scale [76]. Validation standards require community consensus on evidentiary requirements for causal claims from observational data, analogous to established frameworks for randomised trials but adapted to healthcare’s unique constraints. Regulatory pathways must evolve to evaluate systems that make interventional predictions and counterfactual reasoning, extending beyond traditional accuracy-based assessment to mechanistic validity. Clinical trials demonstrating impact on patient outcomes rather than surrogate metrics prove essential to justify substantial implementation costs. Training the next generation of clinician-scientists fluent in both causal inference theory and clinical medicine will determine whether these methods reach their potential.
The transition from correlation-based prediction to mechanism-based medicine will define which artificial intelligence systems earn clinical trust and achieve sustainable deployment. This transformation requires balancing theoretical ambition with empirical humility, acknowledging that we construct pragmatic causal approximations that support clinical inferences at chosen levels of abstraction rather than discovering absolute causal truth in biological systems of extraordinary complexity. It demands computational sophistication matched with clinical interpretability, ensuring that mechanistic insights align with established biomedical knowledge and can be verified by domain experts. Most critically, it necessitates a transformative vision tempered by uncompromising validation standards, where novel causal discoveries undergo rigorous multi-modal triangulation before informing high-stakes clinical decisions. The methodological foundations are established through theoretical advances connecting structural causal models to differentiable graph architectures, empirical demonstrations across diverse healthcare applications, and emerging frameworks for validation under observational constraints [42, 13, 14]. The path forward demands not only technical innovation, but sustained collaboration between computational scientists developing causal inference methods, clinicians defining requirements and validation criteria, biomedical researchers providing mechanistic knowledge, and policymakers establishing appropriate oversight. Multi-modal biomedical data accumulation, regulatory demands for explainable artificial intelligence, and documented failures of correlation-based models under distribution shift together create both opportunity and urgency. Translation to clinical practice would enable healthcare systems to identify causal mechanisms rather than mere associations, predict interventional outcomes rather than observational patterns, and generate patient-specific counterfactuals that enable truly personalised medicine. This vision warrants continued investment provided that we maintain rigorous standards for genuine causal inference in complex biological systems, distinguish mechanistic understanding from pattern recognition, and prioritise patient outcomes over algorithmic sophistication.
Acknowledgements
MM is supported by the Rhodes Trust and the EPSRC CDT Health Data Science. TZ is supported by the Royal Academy of Engineering.
[…]
Author contributions
[…]
Competing interests We declare no competing interests.
[…]
Supplementary information
Supplementary material is contained in a separate file.